OpenStudy (anonymous):

Tell me how to solve the Experimental Probability because im having trouble solving these types of questions. I need this fast.

OpenStudy (phi):

you will have to post a specific problem.

OpenStudy (anonymous):

Tyler has made 65 out of his last 150 free throw attempts. Estimate the probability that he will make his next free throw.

OpenStudy (anonymous):

Tell me how to solve

OpenStudy (phi):

form the ratio (or fraction, which means the same thing) of \[ \frac{number\ of\ successes}{total\ number} \]

OpenStudy (anonymous):

Also tell me how to solve this A movie theater sells popcorn in small, medium, large and jumbo sizes. The customers of the first show purchase 4 small, 20 medium, 40 large, and 16 jumbo containers of popcorn. Find the probability shown below.

OpenStudy (phi):

did you do the first one yet?

OpenStudy (anonymous):

Yes I got 0.433333

OpenStudy (anonymous):

So he has a 43% chance?????????

OpenStudy (phi):

yes, if you round to the nearest whole number

OpenStudy (anonymous):

Thanks and if you can try solving the second question I gave you

OpenStudy (phi):

It asks Find the probability shown below.

OpenStudy (anonymous):

Find the probability shown below.

OpenStudy (phi):

yes, and you have not posted that

OpenStudy (anonymous):

Sorry Im a starter to this

OpenStudy (phi):

Did you find the rest of the question?

OpenStudy (anonymous):

P(mediam)

OpenStudy (phi):

I assume you mean Probability of medium size. It is the same idea: \[ \frac{number\ of\ medium}{total\ number} \]

OpenStudy (anonymous):

Thank you very much!!!

OpenStudy (phi):

They did not tell you the total, you have to add up the number of each size to get it.

OpenStudy (anonymous):

OK I will use this equation for future questions Thank You for the advice Sorry I dint tell you what part of the question was

OpenStudy (phi):

yw

OpenStudy (phi):

Did you find the answer? you should get 20/80 = 0.25 or 25%

OpenStudy (anonymous):

4+20+40+16=80 mediam= 20 20/80=0.25?????????

OpenStudy (phi):

yes